OpenStudy (dollbeautyx):

Let y represent the total cost of publishing a book (in dollars). Let x represent the number of copies of the book printed. Suppose that x and y are related by the equation y=1100+15x

jhonyy9 (jhonyy9):

so and what is the question please ?

OpenStudy (dollbeautyx):

what is the change in total cost for each book printed?

OpenStudy (dollbeautyx):

what is the cost to got started (before any books are printed)?

OpenStudy (dollbeautyx):

get*

jhonyy9 (jhonyy9):

from what you have wrote above result that you need calcule the value of y for x=0 - yes ?

jhonyy9 (jhonyy9):

do you can calcule it so in this way ?

OpenStudy (dollbeautyx):

yes

OpenStudy (dollbeautyx):

y= x=

jhonyy9 (jhonyy9):

ok this mean that so you will get the right answer to this first question

jhonyy9 (jhonyy9):

what mean this y= and x= ???

OpenStudy (dollbeautyx):

this whole equation is is just confusing, I have no idea how to solve it. Can you walk me through the steps?

jhonyy9 (jhonyy9):

y=1100 +15x so how you have wrote in the text of your exercise y mean the cost of x printed books - yes ?

OpenStudy (dollbeautyx):

yes

jhonyy9 (jhonyy9):

so for the start when the printed booke number is zero so just you need substituting the zero in place of x inside this equation of y and make the calcule so will get the started necessary cost

jhonyy9 (jhonyy9):

y=1100 +15x y=1100 +15*0 y = ?

OpenStudy (dollbeautyx):

1100?

jhonyy9 (jhonyy9):

exactly - so was hard ?

OpenStudy (dollbeautyx):

not at all, thanks.

jhonyy9 (jhonyy9):

was my pleasure good luck bye bye